OpenStudy (anonymous):

deffarentiate plz ((5x)^sqrtx)^x

OpenStudy (anonymous):

can u use the equation i can't mean the problem

OpenStudy (lalaly):

\[(5x ^{\sqrt{x}})^{x}\]

OpenStudy (anonymous):

yes lalaly

OpenStudy (anonymous):

\[5^x \left(x^{\sqrt{x}}\right)^x \text{Log}[5]+5^x \left(x^{\sqrt{x}}\right)^x \left(x \left(\frac{1}{\sqrt{x}}+\frac{\text{Log}[x]}{2 \sqrt{x}}\right)+\text{Log}\left[x^{\sqrt{x}}\right]\right) \]

OpenStudy (anonymous):

let y=(\[5x^\sqrt{x}\])^x logy=\[xlog5x^\sqrt{x}\] logy=\[x^\sqrt{x}\log5x\] 5x(xx√)xLog[5]+5x(xx√)x(x(1x√+Log[x]2x√)+Log[xx√])ans then u=\[x^\sqrt{x}\] logu=\[\sqrt{x}logx\] i/u du/dx=\[\sqrt{x}/x+1/2\sqrt{x}logx\] \[du/dx=x^\sqrt{x}\left[ \sqrt{x}/x+1/2\sqrt{xlogx} \right]\] then put the value du/dx=\[du/dx=x^\sqrt{x}\left[ \sqrt{x}/x+1/2\sqrt{xlogx} \right]\]

OpenStudy (anonymous):

aroub its not a simple equation here u need to use calculus